OpenStudy (anonymous):

fxdx=1/2 f(x)= 1/72x [0,12] find the median m

OpenStudy (anonymous):

can you post the question?

OpenStudy (anonymous):

Let x be the continuous random variable over [a,b] with probability density function f. Then the median of the x-values is that number m such that f(x)dx=1/2 f(x)=1/72x,[0,12] find the median m

OpenStudy (anonymous):

before f(x)dx it has that big S with m at the top and a at the bottom

OpenStudy (anonymous):

suddenly the probability got hard hmmm

OpenStudy (anonymous):

i know how to do it normally but not to find m

OpenStudy (anonymous):

is it \[f(x)=\frac{1}{72}x\]?

OpenStudy (anonymous):

yes thats how f(x) looks

OpenStudy (anonymous):

but im pretty sure you take the antiderivative of it

OpenStudy (anonymous):

i guess you have to solve \[\int_0^b\frac{xdx}{72}=\frac{1}{2}\] for \(b\)

OpenStudy (anonymous):

anti derivative is easy enough, it is \(\frac{x^2}{144}\)

OpenStudy (anonymous):

and since if you evaluate it as zero you got zero, you are left with solving \[\frac{x^2}{144}=\frac{1}{2}\]

OpenStudy (anonymous):

so x^2=72? @satellite73

OpenStudy (anonymous):

where does m come into play?

OpenStudy (anonymous):

that is all \[x=\sqrt{72}\]

OpenStudy (anonymous):

thats the median?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sweet thank you!

OpenStudy (anonymous):

yw